A point is moving along the X axis with a constant speed. it is initially located at +2.0 m and 180 seconds later it is located at -5.0 m. what is the average velocity of the point


1
Expert's answer
2021-03-08T08:28:35-0500
"v=\frac{x_2-x_1}{t}\\v=\frac{-5-2}{180}=-0.039\frac{m}{s}"